OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the first and third equations, then the second and third equations. x – y + 2z = –2 2x + 2y + z = 7 3x + 3y – z = 3 2x + 2y = 7 3x + 3y = 3 2x + 2y = 7 5x + 5y = 10 7x + 5y = 4 3x + 3y = 3 7x + 5y = 4 5x + 5y = 10

OpenStudy (anonymous):

@ghazi , @experimentX ?

OpenStudy (ghazi):

what's your point in it ? do you wants us to check your answer?

OpenStudy (anonymous):

first multiply the 3rd equation by 2 & add it to the 1st equation to eliminate z. What do you get?

OpenStudy (ghazi):

@yummydum i think you've already done everything

OpenStudy (ghazi):

wait there is a mistake

OpenStudy (anonymous):

@ghazi this is a question with its options not a worked out problem

OpenStudy (ghazi):

okay ....

OpenStudy (ghazi):

add equation 2 and 3

OpenStudy (ghazi):

what you got?

OpenStudy (anonymous):

@rajathsbhat 7x+5y=4 and then id get 5x+5y=10

OpenStudy (anonymous):

so its D :) ..THNAKYOU :)

OpenStudy (ghazi):

just add these 2x+2x+z=7 and 3x+3y-z=3

OpenStudy (anonymous):

there you go :)

OpenStudy (ghazi):

that's correct